# Discrete conformal structures on surfaces with boundary (II)—Rigidity and Existence

### v1

@@ 6.1. Admissible space of mixed discrete conformal structure $\mathrm{II}$
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,0)$. Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$. Then
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,0)$. Then $u_{j}+u_{k}>\log\frac{2}{\xi_{jk}},\ \xi_{jk}>0$. Moreover, $\xi_{ij}>0$$u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$ and $\xi_{ik}>0,\ u_{i}+u_{k}>\log\frac{1}{\xi_{ik}}$. Then
$\mathcal{U}^{(3)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}|u_{j}+u_{k}>\log\frac{2}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},\ u_{i}+u_{k}>\log\frac{1}{\xi_{ik}}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1. Moreover, $\xi_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$. Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$ and $u_{i}+u_{k}>\log(-\xi_{ik})$ for $\xi_{ik}<0$. Then
(50) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ e^{u_{i}+u_{k}}>-\xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\in\mathbb{R}$.
The admissible space (50) can be described explicitly as follows:
@@
For $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\geq 0$, then
(51) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}\}.$
@@
For $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}<0$, then
(52) $\mathcal{U}^{(4)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}>-1$. Then
(53) $\mathcal{U}^{(5)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}_{>0}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}\},$
where $\xi_{jk}>0,\ \xi_{ij}\geq 0,\ \xi_{ik}>-1$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}>0$$u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})$. Then
$\mathcal{U}^{(6)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}|$ $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},$
$u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}\in\mathbb{R}$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,-1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$ in Subsection 3.1. Moreover, $\xi_{ij}>0$$(u_{j}-u_{i})\in\mathbb{R}$. Similarly, $\xi_{ik}>0$ and $u_{i}+u_{k}<\log\xi_{ik}$. Then
$\mathcal{U}^{(7)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}|u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{k}<\log\xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>0$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,-1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$. Similarly, $\xi_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\ \xi_{ik}$. Then
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,-1)$. Then $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}}$ and $\xi_{jk}>0$. Moreover, $\xi_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\xi_{ij}}$. Similarly, $\xi_{ik}>1$ and $-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}$. Then
$\mathcal{U}^{(9)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}|$ $u_{j}+u_{k}>\log\frac{1}{\xi_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\xi_{ij}},$
$-\mathrm{arccosh}\ \xi_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\ \xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ij}>0,\ \xi_{ik}>1$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$ in Subsection 3.1. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $u_{i}+u_{j}>\log(-\xi_{ij})$ for $\xi_{ij}<0$. Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\xi_{ik}\geq 0$ and $u_{i}+u_{k}>\log(-\xi_{ik})$ for $\xi_{ik}<0$. Then
(54) $\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ e^{u_{i}+u_{j}}>-\xi_{ij},$
$e^{u_{i}+u_{k}}>-\xi_{ik}\},$
where $\xi_{jk}>1,\ \xi_{ij}\in\mathbb{R},\ \xi_{ik}\in\mathbb{R}$.
The admissible space (54) can also be described explicitly as follows:
@@
For $\xi_{jk}>1,\ \xi_{ij}\geq 0,\ \xi_{ik}\geq 0$, then
$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}\}.$
@@
For $\xi_{jk}>1,\ \xi_{ij}\geq 0,\ \xi_{ik}<0$, then
$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{k}>\log(-\xi_{ik})\}.$
@@
For $\xi_{jk}>1,\ \xi_{ij}<0,\ \xi_{ik}\geq 0$, then
$\mathcal{U}^{(10)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},\ u_{i}+u_{j}>\log(-\xi_{ij})\},$
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}>-1$$(u_{i}+u_{j})\in\mathbb{R}$ and $\xi_{ik}>-1,\ (u_{i}+u_{k})\in\mathbb{R}$. Then
(55) $\mathcal{U}^{(11)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}\},$
where $\xi_{jk}>1,\ \xi_{ij}>-1,\ \xi_{ik}>-1$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,1)$. Then $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk}$ and $\xi_{jk}>1$. Moreover, $\xi_{ij}\in\mathbb{R}$$u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij})$ and $\xi_{ik}\in\mathbb{R},\ u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})$. Then
$\mathcal{U}^{(12)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>-\mathrm{arccosh}\ \xi_{jk},u_{i}+u_{j}>\mathrm{arcsinh}\ (-\xi_{ij}),$
$u_{i}+u_{k}>\mathrm{arcsinh}\ (-\xi_{ik})\},$
where $\xi_{jk}>1,\ \xi_{ij}\in\mathbb{R},\ \xi_{ik}\in\mathbb{R}$.
@@
Set $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,-1)$. Then $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk})$ and $\xi_{jk}>0$ in Subsection 3.1. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\xi_{ij}\geq 0$ and $u_{i}+u_{j}>\log(-\xi_{ij})$ for $\xi_{ij}<0$. Similarly, $\xi_{ik}>0$ and $u_{i}+u_{k}<\log\xi_{ik}$. If $\xi_{ij}<0$, then $\log(-\xi_{ij})+\mathrm{arcsinh}\ (-\xi_{jk})<u_{i}+2u_{j}+u_{k}<2u_{j}+\log\xi_{ik}<\log\xi_{ik}$, which implies $(\xi_{ik}-\xi_{ij})(\xi_{ik}+\xi_{ij})>2\xi_{jk}\xi_{ij}\xi_{ik}$. Thus we assume $\xi_{ik}+\xi_{ij}>0$ for simplicity. Then
$\mathcal{U}^{(13)}_{ijk}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ e^{u_{i}+u_{j}}>-\xi_{ij},$
$u_{i}+u_{k}<\log\xi_{ik}\},$
where $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}\geq 0$ or $\xi_{ij}<0$$\xi_{ij}+\xi_{ik}>0$.
The admissible space $\mathcal{U}^{(13)}_{ijk}(\xi)$ can be described explicitly as follows:
@@
For $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}\geq 0$, then
$\mathcal{U}_{ijk}^{(13)}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),u_{i}+u_{k}<\log\xi_{ik}\}.$
@@
For $\xi_{jk}>0,\ \xi_{ik}>0,\ \xi_{ij}<0$ and $\xi_{ij}+\xi_{ik}>0$, then
$\mathcal{U}_{ijk}^{(13)}(\xi)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}|$ $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\xi_{jk}),\ u_{i}+u_{k}<\log\xi_{ik},$
$u_{i}+u_{j}>\log(-\xi_{ij})\}.$

### v3

@@ 8.1. Admissible space of the mixed discrete conformal structure I
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,0)$. It is known that $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}\geq 0$. Consequently, the admissible space of $u$ is defined as
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,0)$. It is known that $u_{j}+u_{k}>\log\frac{2}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $\eta_{ij}>0$ with $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$, and $\eta_{ik}>0$ with $u_{i}+u_{k}>\log\frac{1}{\eta_{ik}}$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(3)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}\mid u_{j}+u_{k}>\log\frac{2}{\eta_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\eta_{ij}},\ u_{i}+u_{k}>\log\frac{1}{\eta_{ik}}\},$
where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,1)$. In Subsection 3.1, it follows that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $\eta_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$. Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}\geq 0$, and $u_{i}+u_{k}>\log(-\eta_{ik})$ for $\eta_{ik}<0$. Consequently, the admissible space of $u$ is defined as
(57) $\mathcal{U}^{(4)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ e^{u_{i}+u_{k}}>-\eta_{ik}\},$
where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}\in\mathbb{R}$.
The admissible space (57) can be explicitly characterized as follows:
@@
For $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}\geq 0$, the admissible space is given by
(58) $\mathcal{U}^{(4)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}\}.$
@@
For $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}<0$, the admissible space is given by
(59) $\mathcal{U}^{(4)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{k}>\log(-\eta_{ik})\}.$
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,1)$. It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}>-1$. Consequently, the admissible space of $u$ is defined as
(60) $\mathcal{U}^{(5)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}_{>0}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}\},$
where $\eta_{jk}>0$, $\eta_{ij}\geq 0$, and $\eta_{ik}>-1$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,1)$. It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $\eta_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$. Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(6)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}\mid$ $\ u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\eta_{ij}},$
$\ u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})\},$
where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}\in\mathbb{R}$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,0,-1)$. In Subsection 3.1, it follows that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $\eta_{ij}>0$ and $(u_{j}-u_{i})\in\mathbb{R}$. Similarly, $\eta_{ik}>0$ and $u_{i}+u_{k}<\log\eta_{ik}$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(7)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}^{2}\times\mathbb{R}_{<0}\mid u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{k}<\log\eta_{ik}\},$
where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>0$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,0,-1)$. It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$. Similarly, $\eta_{ik}\in\mathbb{R}$ and $u_{i}+u_{k}<\mathrm{arcsinh}\,\eta_{ik}$. Consequently, the admissible space of $u$ is defined as
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,0,-1)$. It is known that $u_{j}+u_{k}>\log\frac{1}{\eta_{jk}}$ and $\eta_{jk}>0$. Moreover, $\eta_{ij}>0$ and $u_{i}+u_{j}>\log\frac{1}{\eta_{ij}}$. Similarly, $\eta_{ik}>1$ and $-\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(9)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}\times\mathbb{R}_{<0}\mid$ $\ u_{j}+u_{k}>\log\frac{1}{\eta_{jk}},\ u_{i}+u_{j}>\log\frac{1}{\eta_{ij}},$
$\ -\mathrm{arccosh}\,\eta_{ik}<u_{i}+u_{k}<\mathrm{arccosh}\,\eta_{ik}\},$
where $\eta_{jk}>0$, $\eta_{ij}>0$, and $\eta_{ik}>1$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,1)$. In Subsection 3.1, it follows that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $u_{i}+u_{j}>\log(-\eta_{ij})$ for $\eta_{ij}<0$. Similarly, $(u_{i}+u_{k})\in\mathbb{R}$ for $\eta_{ik}\geq 0$, and $u_{i}+u_{k}>\log(-\eta_{ik})$ for $\eta_{ik}<0$. Consequently, the admissible space of $u$ is defined as
(61) $\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ e^{u_{i}+u_{j}}>-\eta_{ij},$
$\ e^{u_{i}+u_{k}}>-\eta_{ik}\},$
where $\eta_{jk}>1$, $\eta_{ij}\in\mathbb{R}$, and $\eta_{ik}\in\mathbb{R}$.
The admissible space (61) can be explicitly characterized as follows:
@@
For $\eta_{jk}>1$, $\eta_{ij}\geq 0$, and $\eta_{ik}\geq 0$, the admissible space is given by
$\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}\}.$
@@
For $\eta_{jk}>1$, $\eta_{ij}\geq 0$, and $\eta_{ik}<0$, the admissible space is given by
$\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{k}>\log(-\eta_{ik})\}.$
@@
For $\eta_{jk}>1$, $\eta_{ij}<0$, and $\eta_{ik}\geq 0$, the admissible space is given by
$\mathcal{U}^{(10)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{j}>\log(-\eta_{ij})\},$
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(1,1,1)$. It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$. Moreover, $\eta_{ij}>-1$ with $(u_{i}+u_{j})\in\mathbb{R}$, and $\eta_{ik}>-1$ with $(u_{i}+u_{k})\in\mathbb{R}$. Consequently, the admissible space of $u$ is defined as
(62) $\mathcal{U}^{(11)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}\},$
where $\eta_{jk}>1$, $\eta_{ij}>-1$, and $\eta_{ik}>-1$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(-1,1,1)$. It is known that $u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk}$ and $\eta_{jk}>1$. Moreover, $\eta_{ij}\in\mathbb{R}$ with $u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij})$, and $\eta_{ik}\in\mathbb{R}$ with $u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(12)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}_{>0}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>-\mathrm{arccosh}\,\eta_{jk},\ u_{i}+u_{j}>\mathrm{arcsinh}\,(-\eta_{ij}),$
$\ u_{i}+u_{k}>\mathrm{arcsinh}\,(-\eta_{ik})\},$
where $\eta_{jk}>1$, $\eta_{ij}\in\mathbb{R}$, and $\eta_{ik}\in\mathbb{R}$.
@@
Assume $(\alpha_{i},\alpha_{j},\alpha_{k})=(0,1,-1)$. In Subsection 3.1, it follows that $u_{j}+u_{k}>\mathrm{arcsinh}\ (-\eta_{jk})$ and $\eta_{jk}>0$. Moreover, $(u_{i}+u_{j})\in\mathbb{R}$ for $\eta_{ij}\geq 0$, and $u_{i}+u_{j}>\log(-\eta_{ij})$ for $\eta_{ij}<0$. Similarly, $\eta_{ik}>0$ and $u_{i}+u_{k}<\log\eta_{ik}$. When $\eta_{ij}<0$, the inequalities $\log(-\eta_{ij})+\mathrm{arcsinh}\,(-\eta_{jk})<u_{i}+2u_{j}+u_{k}<2u_{j}+\log\eta_{ik}<\log\eta_{ik}$ implies $(\eta_{ik}-\eta_{ij})(\eta_{ik}+\eta_{ij})>2\eta_{jk}\eta_{ij}\eta_{ik}$. For simplicity, we assume $\eta_{ik}+\eta_{ij}>0$. Consequently, the admissible space of $u$ is defined as
$\mathcal{U}^{(13)}_{ijk}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ e^{u_{i}+u_{j}}>-\eta_{ij},$
$\ u_{i}+u_{k}<\log\eta_{ik}\},$
where $\eta_{jk}>0$, $\eta_{ik}>0$, and $\eta_{ij}$ satisfies either $\eta_{ij}\geq 0$ or $\eta_{ij}<0$ with $\eta_{ij}+\eta_{ik}>0$.
The admissible space $\mathcal{U}^{(13)}_{ijk}(\eta)$ can be explicitly characterized as follows:
@@
For $\eta_{jk}>0$, $\eta_{ik}>0$, and $\eta_{ij}\geq 0$, the admissible space is given by
$\mathcal{U}_{ijk}^{(13)}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),u_{i}+u_{k}<\log\eta_{ik}\}.$
@@
For $\eta_{jk}>0$, $\eta_{ik}>0$, $\eta_{ij}<0$, and $\eta_{ij}+\eta_{ik}>0$, the admissible space is given by
$\mathcal{U}_{ijk}^{(13)}(\eta)=\{(u_{i},u_{j},u_{k})\in\mathbb{R}\times\mathbb{R}^{2}_{<0}\mid$ $\ u_{j}+u_{k}>\mathrm{arcsinh}\,(-\eta_{jk}),\ u_{i}+u_{k}<\log\eta_{ik},$
$\ u_{i}+u_{j}>\log(-\eta_{ij})\}.$